Sally drives 29.0 min at 59.0 km/h, 35.0 min at 45.0 km/h and 45.0 min at 69.0 and she spends 6.0 min eating lunch and buying gas, determine the average speed
1
Expert's answer
2012-12-10T10:47:26-0500
V - an average speed
S - a total distance
t - a total time

V = S/t
S = V1*t1 + V2*t2 + V3*t3
t = t1 + t2 + t3 + t4
V = (29*59 + 35*45 + 45*69) / (29 +35 + 45 + 6) = (1711 + 1575 + 3105)/115 =
6391/115 = 55.6 km/h

Answer: the average speed is 55.6 km/h